A monochromatic light source of intensity 5 mW emits 8 × 10^15 photons per second. This light ejects photoelectrons from a metal surface. The stopping potential for this setup is 2.0 V. Calculate the work function of the metal.

Grade:upto college level

1 Answers

Aditi Chauhan
9 years ago
Sol. W base 0 hv – ev base 0 = 5 * 10^-3/8 * 10^15 – 1.6 * 10^-19 * 2 (Given V base 0 = 2V, No. of photon = 8 * 10^15, Power = 5 mW) = 6.25 * 10^-19 – 3.2 * 10^-19 = 3.05 * 10^-19 J = 3.05 * 10^-19/1.6 * 10^-19 = 1.906 eV.